# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v2

@@ 7.2.6 Stability of PPFPL
Fig. 8: Stability of PPFPL under MNIST, FMNIST, and CIFAR10.
It is observed from Fig. 4 and 5 that the performance of PPFPL fluctuates. Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments. To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of participants being malicious. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). Each experiment is repeated 10 times. The utilization of boxplots in Fig. 8 provides a clear visualization of the stability of PPFPL. The results show that PPFPL’s performance fluctuates less and remains stable against data poisoning attacks. The stability of PPFPL proves its reliability in real-world scenarios.
7.2.7 Scalability of PPFPL
@@
Fig. 9: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
7.2.8 Impact of Local Iterations
@@
Fig. 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
7.2.9 Efficiency Evaluation
@@
Fig. 11: Communication parameters. (a) The number of parameters submitted by clients under different local models. (b) The number of parameters submitted by client using CNN model in one round under different scheme.
To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients in PPFPL, as shown in Fig. 11. We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same parameters. Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture. This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, not the model architecture. Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
Table 3: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
@@
In addition, ciphertext operations are the most important factor affecting the efficiency of our framework. We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations. Note that PBFL is a federated learning scheme based on CKKS, while ShieldFL is built upon the two-door HE. For each client, we use the time spent on encryption and decryption for each iteration as a metric. As shown in Table 3, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL. This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
7.3 Complexity Analysis
Table 4: Computation Overload and Communication Overload of client.
@@ 7.3 Complexity Analysis
We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15]. The results are shown in Table 4. The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption. Formally, the computational overhead is expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ denotes the prototype parameters, $T_{ch}$ denotes the time overload of encryption, $T_{tr}$ denotes the time overhead of local model training, and $T_{pro}$ denotes the time overhead of prototype generation. Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$. Other similar schemes [16][18][15] adopt homomorphic encryption on gradients, their encryption overhead is $\mathcal{O}(gT_{ch})$, where $g$ denotes the gradient parameters. Since $g>p$, the overload satisfies $\mathcal{O}(gT_{ch})>\mathcal{O}(pT_{ch})$. Hence, the relationship $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ holds. Consequently, PPFPL imposes a reasonable computational overhead on clients.
@@ 8 Concludes
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grants 62272256 and 62202250, the Major Program of Shandong Provincial Natural Science Foundation for the Fundamental Research under Grant ZR2022ZD03, the National Science Foundation of Shandong Province under Grant ZR2021QF079, the Talent Cultivation Promotion Program of Computer Science and Technology in Qilu University of Technology (Shandong Academy of Sciences) under Grant 2023PY059, the Pilot Project for Integrated Innovation of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences) under Grant 2022XD001, and the Colleges and Universities 20 Terms Foundation of Jinan City under Grant 202228093.
CRediT authorship contribution statement
Hongliang Zhang: Writing – original draft, Validation, Software, Methodology, Conceptualization. Jiguo Yu: Writing – review & editing, Supervision. Fenghua Xu: Software, Validation, Writing – review & editing. Chunqiang Hu: Writing – review & editing. Yongzhao Zhang: Writing – review & editing, Supervision. Xiaofen Wang: Software, Validation. Zhongyuan Yu: Validation. Xiaosong Zhang: Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.

### v4

@@ VII-B6 Stability of PPFPL
Figure 8: Stability of PPFPL under MNIST, FMNIST, and CIFAR10.
It is observed from Fig. 4 and 5 that the performance of PPFPL exhibits some fluctuations. Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments. To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). Each experiment is repeated 10 times. The utilization of boxplots in Fig. 8 provides a clear visualization of the stability of PPFPL. The results show that PPFPL’s performance fluctuates less and remains stable against data poisoning attacks. The stability of PPFPL proves its reliability in real-world scenarios.
VII-B7 Scalability of PPFPL
@@
Figure 9: Scalability of PPFPL against data poisoning attacks over the MNIST, FMNIST, and CIFAR10.
VII-B8 Impact of Local Iterations
@@
Figure 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
VII-B9 Efficiency Evaluation
@@
Figure 11: Communication parameters. (a) The number of parameters submitted by clients under different local models. (b) The number of parameters submitted by client using CNN model in one round under different scheme.
To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as shown in Fig. 11. We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same amount of parameters. Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture. This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, rather than the model architecture. Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
TABLE III: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
@@
In addition, ciphertext operations are the primary factor affecting the efficiency of our framework. We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations. Note that PBFL is a federated learning scheme based on CKKS, while ShieldFL is built upon the two-door HE. For each client, we use the time spent on encryption and decryption for each iteration as a metric. As shown in Table III, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL. This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
VII-C Complexity Analysis
TABLE IV: Computation Overload and Communication Overload of client.
@@ VII-C Complexity Analysis
We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15]. The results are shown in Table IV. The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption. Formally, the computational overhead is expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ denotes the prototype parameters, $T_{ch}$ denotes the time overload of encryption, $T_{tr}$ denotes the time overhead of local model training, and $T_{pro}$ denotes the time overhead of prototype generation. Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$. Other similar schemes [16][18][15] adopt homomorphic encryption on gradients, their encryption overhead is $\mathcal{O}(gT_{ch})$, where $g$ denotes the gradient parameters. Since $g>p$, the overload satisfies $\mathcal{O}(gT_{ch})>\mathcal{O}(pT_{ch})$. Hence, the relationship $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ holds. Consequently, PPFPL imposes a reasonable computational overhead on clients.